What is the mass in grams of 34.1 amu?

Enter your answer in scientific notation.
1
Expert's answer
2019-10-14T14:31:33-0400

Solution.

"1 \ amu = 1.66 \times 10^{-24} \ g"

Then,

"34.1 \ amu = 1.66 \times 10^{-24} \times 34.1 = 5.66 \times 10^{-23} \ g"

Answer:

"34.1 \ amu = 5.66 \times 10^{-23} \ g"